OpenStudy (anonymous):

Jake and Seth put some money into their money boxes every week. The amount of money (y), in dollars, in their money boxes after a certain amount of time (x), in weeks, is shown by the equations below: Jake: y = 5x + 85 Seth: y = 20x + 40 After how many weeks will Jake and Seth have the same amount of money in their money boxes and what will that amount be? A. 3 weeks, $100 B. 3 weeks, $285 C. 6 weeks, $150 D. 100 weeks, $3

OpenStudy (anonymous):

Solve the following for x: 5 x + 85 = 20 x + 40 x = 3

OpenStudy (anonymous):

i still need help!!!

OpenStudy (anonymous):

When Jake and Seth have the same amount of money, then: 5 x + 85 = 20 x + 40 x is the amount of time and in the case we are concerned when, 5 x + 85 = 20 x + 40